Evaluate the ∫(t³+3t)/(t²+1)dt


1
Expert's answer
2021-03-01T15:21:23-0500


Let "I=\int \dfrac{t^3+3t}{t^2+1} dt"

First performing long division, we get

"I=\int (t+\dfrac{2t}{t^2+1} )dt"

On integrating

"I=\dfrac{t^2}2+\log|t^2+1|+C"